OpenStudy (anonymous):

Which of the following side lengths do NOT form a right triangle? A. 5 cm, 12 cm and 13 cm B. 8 cm, 15 cm and 17 cm C. 9 in, 40 in and 41 in D. 20 in, 21 in and 28 in

HanAkoSolo (jamierox4ev3r):

follow this guideline: a+b>c b+c>a a+c>b

HanAkoSolo (jamierox4ev3r):

but for a right triangle, you would test the pythagorean method

HanAkoSolo (jamierox4ev3r):

a^2+b^2=c^2 This is the Pythagorean theorem. Can you figure that out for me?

OpenStudy (anonymous):

should i try out all the angles ? @Jamierox4ev3r

HanAkoSolo (jamierox4ev3r):

it would be good so that you can make sure your answer is correct

HanAkoSolo (jamierox4ev3r):

mention my name when you have the answer

OpenStudy (anonymous):

D. Does 20^2+21^2=28^2 400 + 441 = 841 and 28^2=784 No, not a right triangle @Jamierox4ev3r

HanAkoSolo (jamierox4ev3r):

in this case, after you proved that, A. is the correct answer because of the following: 5^2+12^2=169 the square root of 169=13 Therefore, 5^2+12^2=13^2

HanAkoSolo (jamierox4ev3r):

Does that make sense?

OpenStudy (anonymous):

not really @Jamierox4ev3r

HanAkoSolo (jamierox4ev3r):



HanAkoSolo (jamierox4ev3r):

5^2=25 12^2=144 144+25=169 169=13^2 Therefore, 12^2+5^2=13^2

HanAkoSolo (jamierox4ev3r):

all of this proves that the above triangle is a right triangle indeed

OpenStudy (anonymous):

so i can but A for the answer @Jamierox4ev3r

OpenStudy (anonymous):

A. Does 5^2+12^2=13^2 25 + 144 = 169 and 13^2=169 Yes, right triangle B. Does 8^2+15^2=17^2 64 + 225 = 289 and 17^2=28 Yes, right triangle C. Does 9^2+40^2=41^2 81 + 1600 = 1681 and 41^2=1681 Yes, right triangle D. Does 20^2+21^2=28^2 400 + 441 = 841 and 28^2=784 No, not a right triangle @Jamierox4ev3r

HanAkoSolo (jamierox4ev3r):

oh idk what i was thinking for some if you are talking which is not a right triangle, the answer would be D. lol sorry

HanAkoSolo (jamierox4ev3r):

im slacking off srry :)) and thank you for pointing out my mistake

OpenStudy (anonymous):

thanks yay ! (: @Jamierox4ev3r

HanAkoSolo (jamierox4ev3r):

anytime just keep up the great work and DFTBA (don't forget to be awesome)!